OpenStudy (anonymous):

what is the simplified form of 4z^2-4z-15/2z^2+z-15 ?

OpenStudy (mertsj):

\[\frac{4z^2-4z-15}{2z^2+z-15}=\frac{(2z-5)(2z+3)}{(2z-5)(z+3)}=\frac{2z+3}{z+3}\]